### v1

@@ Abstract
[style=chinese]
1]organization=Noakhali Science and Technology University, addressline=Sonapur, Noakhali, postcode=3814, country=Bangladesh
[style=chinese]
2]organization=University of Southern Mississippi, addressline=Hattiesburg, Mississippi, citysep=, postcode=394901 , country=USA
3]organization=Universiti Malaysia Pahang, addressline=Pekan, Pahang, citysep=, postcode=26600 , country=Malaysia
[style=chinese]
4]organization=King Saud University, addressline=Riyadh, citysep=, postcode=11451, country=Saudi Arabia
[style=chinese]
5]organization=Khulna University, addressline=Khulna, citysep=, postcode=9208 , country=Bangladesh
[style=chinese]
6]organization=Cybersecurity and FinTech Computer Science Department, City, University of London, addressline=EC1V 0HB, citysep=, postcode= , country=UK
[…]
[cor1]Anupam Kumar Bairagi, Sujit Biswas
However, the MpoxSLDNet model achieved a validation accuracy of 94.56%, compared to 86.25%, 84.38%, and 67.19% for VGG16, DenseNet121, and ResNet50, respectively. In addition, an average precision, recall, F1 score of 0.94 respectively, and AUC of 0.94 was recorded for our proposed model.
keywords:
[…]
However, the MpoxSLDNet model achieved a validation accuracy of 94.56%, compared to 86.25%, 84.38%, and 67.19% for VGG16, DenseNet121, and ResNet50, respectively. In addition, an average precision, recall, F1 score of 0.94 respectively, and AUC of 0.94 was recorded for our proposed model.
1 Introduction
@@
Figure 1 provides an introductory overview of the efforts to enhance monkeypox lesion detection. It begins by elucidating the characteristics of the Monkeypox virus including its origins, classification, symptoms, and modes of transmission. Monkeypox virus (MPXV) is a zoonotic virus that is predominantly found in distant Central and West African regions R1; R2. It was first identified in a monkey in 1959, when outbreaks of a disease resembling pox were observed in monkeys retained for research purposes, in a laboratory in Copenhagen, at a Danish research institute, and thus the name Monkeypox virus R2; R3. It is a type of DNA virus classified within the orthopoxvirus genus and is a member of the family of Poxviridae, is closely linked to the Variola virus, which is the infectious illness that causes monkeypox R4. In 1970, the first recorded instance of monkeypox cases in humans was documented in the Democratic Republic of the Congo, within an intensified campaign to wipe out smallpox. The monkeypox virus falls into two groups: Congo Basin (Central African) and West African. According to a previous report, the Congo Basin clade tends to produce severe disease more often, with case fatality ratios (CFR) of up to 10%. As of right now, the Democratic Republic of Congo reports a 3% Case Fatality Ratio among suspected cases. In the past, the West African clade has been attributed to a lower overall CFR of roughly 1% in an African population who are normally younger. An average incubation period is 12 days for MPXV, yet it might vary from 5 to 21 days. R5. Monkeypox virus signs involve fever, headache, tenderness in the muscles, backaches, swelling of the lymph nodes, chills, as well as exhaustion R6; R7; new2. The monkeypox virus circulates from a human to another human being thru droplets from the nose or throat, interaction with affected bodily fluids or contaminated surfaces, infectious rash, and sexual contact. This virus can also spread through the handling of infected animals R8. There is unfortunately no medication or vaccination specifically against the monkeypox virus. Although supportive care can be given to combat symptoms and undesirable side effects. Antiviral drugs can be used that are used for treating other orthopoxvirus infections R9. A skin lesion test utilizing the technique of electron microscopy or polymerase chain reaction, also known as a PCR test is time-consuming and takes a long time to detect pox R10. The insufficient availability of PCR test kits, as well as the strict requirements for laboratory settings in research facilities, will delay the exact identification of suspects R11. The detection of monkeypox lesions can be challenging due to the similarity that exists between monkeypox lesions and another pox R12; R13, which can lead to misdiagnosis. In addition, personnel in healthcare might have been unfamiliar with a clinical presentation of the lesions of monkeypox R14, which can further complicate accurate diagnosis. Moreover, physical examinations to diagnose monkeypox can increase the risk of transmission to healthcare professionals R15 and other patients. According to R16, isolation is recommended for a patient with a suspected monkeypox infection. To address all these challenges, there is a need for the development of a reliable and accurate diagnostic tool that can quickly and non-invasively detect and classify monkeypox lesions from other poxvirus infections which also can significantly reduce the risk of transmission to healthcare professionals and other patients. Neural networks can help in this situation. With advances in the development involving AI models, many models based on deep learning have been utilized in the identification of various diseases R17; R18; R19; new1. So, in our current research, we have developed a model named “MpoxSLDNet” based on convolutional neural networks (CNNs) to automatically detect monkeypox lesions from digitized pictures of human skin lesions.
@@
From the existing research work regarding monkeypox lesion detection, we can observe, in R1, authors used publicly available datasets. They presented an Android mobile application in which low modified deep learning model MobileNetv2 is used to classify the monkeypox lesion and non-monkeypox lesions with 91.11% accuracy. In R11, authors developed a monkeypox skin lesion dataset, and they employed deep learning networks including VGG16, ResNet50, InceptionV3, and Ensemble. In their study, among the evaluated models, ResNet50 demonstrated the highest overall accuracy, achieving an accuracy of 82.96%. Also VGG16, InceptionV3,Ensemble obtained accuracy of 81.48%, 74.07%,79.26% respectively. Here, the authors presented a web app that used the best-performing model. Apart from this, Authors R21 used publicly available datasets and employed five deep learning models, including DenseNet121, VGG19, MobileNetV2, EfficientNetB3, and Xception combined with the CBAM and Dense layers for categorization. Here, VGG19, DenseNet121, EfficientNetB3, MobileNetV2, and Xception achieved accuracy 71.86%, 78.27%, 81.43%, 74.07%, 83.89% respectively. Despite the significant efforts in the existing research work regarding monkeypox lesion detection, our model achieves superior results on the basis of accuracy, precision, recall, F1-score, and AUC compared to the state-of-the-art models. While the existing thesis work on monkeypox lesion detection has employed several deep-learning models, these models require significant storage space due to their large number of parameters and also require more computational time. In contrast, our proposed CNN model is designed to be both accurate and efficient, with fewer parameters that require a smaller amount of storage as well as computational time. This makes our approach more practical for use in resource-constrained settings, such as remote areas where monkeypox outbreaks are more likely to occur. Our proposed MpoxSLDNet model is a Sequential neural network containing max pooling, convolutional layers, batch normalization, dropout, as well as fully connected layers. It takes input images of size 224x224x3 and learns to classify them as either having monkeypox lesions or not. The feature extraction process in our model involves using a series of convolutional layers of adding depth in addition to reducing spatial dimensions, commencing with max pooling layers to reduce the dimensionality of the feature maps. The extracted features are then flattened and passed through various fully connected layers with batch normalization and dropout regularization to stay away from overfitting. The final layer of output uses a sigmoid activation function to acquire a binary classification result. The contribution of this study to the field of monkeypox lesion detection is listed below:
@@ 2 Literature Review
Monkeypox is a rare and serious viral disease that presents a significant public health risk. The timely detection of monkeypox lesions is essential for effective treatment as well as outbreak prevention. Deep learning-based models have gained prominence in recent years due to their potential use in automating the detection of monkeypox lesions. A recent study conducted by R4 collected monkeypox images from various sources and developed a dataset of monkeypox images that are publicly available. This study established a modified VGG16 model to classify monkeypox patients and non-monkeypox patients. Post-image analysis explanations using a technique called Local Interpretable Model-Agnostic Explanations (LIME) were also used in order to verify outcomes. The modified VGG16 model was evaluated for two distinct studies, including study one (without augmentation) and study two (with augmentation). The achieved accuracies for studies one and two were 97 ± 1.8% and 88 ± 0.8%, respectively. However, this study only considered the modified VGG16 model. Additionally, Another study conducted by R22, proposed a model named MonkeypoxHybridNet based on CNN, which is comprised of VGG19, ResNet50, and InceptionV3. The performance of the model was assessed by evaluating its precision, F1 score, and accuracy against three pre-trained models, including VGG19, ResNet50, and InceptionV3. The achieved accuracy for VGG19, ResNet50, InceptionV3, and MonkeypoxHybridNet were 0.595, 0.705, 0.80, and 0.842, respectively. The proposed model showed significant results, but it has some limitations. The authors only presented accuracy, precision, F1 score, and confusion matrix but did not show the Recall and AUC score for the proposed MonkeypoxHybridNet model. In a subsequent study by R14, a variety of pre-trained models were used to distinguish between chickenpox and monkeypox lesions, including Inception V3, ResNet50, VGG16, VGG19, and AlexNet. They also implemented a CNN model. The chickenpox dataset used in the study was publicly available, while the monkeypox dataset was obtained through web scraping.VGG16 and VGG19 accuracy was also 80%, while ResNet50, AlexNet, and Inception V3 each provided accuracy results of 82.00%, 98.00%, and 89.00%, respectively. The recorded accuracy of the proposed CNN model was 99%. However, the AUC score, which is an important measure of classifier performance, was not calculated in this study. Apart from this, another study by R2 compared the performance of 13 pre-trained models to detect the monkeypox lesion. Universal custom layers had been used for all models, and the performance was analyzed based on four metrics: accuracy, precision, recall, and F1-score. The study identified DL models that exhibited the highest performance and ensembled them via majority voting on their probabilistic outputs to enhance the overall performance. The experiments utilized a dataset that is publicly accessible and the proposed ensemble method provided accuracy, precision, recall, and F1-score of 87.13%, 85.44%, 85.47%, and 85.40%, respectively. However, this study did not propose any lightweight CNN models, and the AUC score was not considered. In a subsequent study conducted by R11, A dataset called the Monkeypox Skin Lesion Dataset (MSLD) was created, which includes images of skin lesions associated with monkeypox, chickenpox, and measles. Three widely used pre-trained models, including InceptionV3, ResNet50, and VGG16, were utilized to identify monkeypox and non-monkeypox patients. Additionally, an ensemble of three models (InceptionV3, ResNet50, and VGG16) was employed in this study. The achieved accuracies were 81.48% (±6.87%), 82.96% (±4.57%), 74.07% (±3.78%), and 79.26% (±1.05%) for VGG16, ResNet50, InceptionV3, and the ensemble method, respectively. Furthermore, a web application was employed as an online monkeypox screening tool. It is known that pre-trained models require more storage space and computational time. In this work, only pre-trained models were utilized, and no lightweight CNN models were established. As for performance evaluation metrics, the AUC score was not considered. Furthermore, a study conducted by R24, used five pre-trained models including VGG16, VGG19, MobileNetV2, ResNet50, and EfficientNetB3. The performance of the model is assessed by measuring its accuracy, recall, precision, and F1 score. Their experimental result demonstrates that among the five pre-trained models MobileNetV2 gives the best classification performance. Achieved accuracy, precision, recall, and F1-score of MobileNetV2 are 98.16%, 0.99, 0.96, and 0.98. Their objective was to select the best-performing model to detect and classify monkeypox. In this study, the authors don’t consider the AUC score and don’t show the multiple training and testing accuracy. They also don’t discuss losses during training. Another study employed by R25, a newly curated image dataset, referred to as the "Monkeypox skin images dataset," was developed, encompassing four classes: monkeypox, chickenpox, measles, and normal images. In this study, five machine learning classifiers, including Logistic Regression, Random Forest, Support Vector Machine, K-Nearest Neighbor, as well as Extreme Gradient Boosting were used for classification. For feature extraction, VGG16, ResNet50, MobileNetV1, Inception V3, Xception. Also, they presented a novel CNN model called MonkeyNet, which is a modified version of DenseNet-201, and evaluated its performance. The performance of the model was assessed using various metrics including accuracy, precision, recall, F1-score, and AUC. Resulted in accuracy for the original dataset and augmented dataset of the proposed model are 91.91% and 98.91%, consecutively. Another study conducted by R21 utilized five deep learning models including VGG19, Xception, DenseNet121, MobileNetV2, and EfficientNetB3.To utilize a more relevant feature map, they integrate the deep learning model with the Convolutional Block Attention Module. This module encompasses mechanisms such as channel attention and spatial attention-based mechanism. Then dense layer, flattened layer, and classification layer are added to each model to classify monkeypox and others. The model’s performance is assessed using accuracy, precision, recall, and F1-score as evaluation metrics. Among the various combinations, Xception-CBAM-Dense yielded the highest performance with a validation accuracy of 83.89%, precision of 90.70%, recall of 89.10%, and F1-score of 90.11%. Authors R32 used deep learning techniques, specifically deep neural networks, for the detection of monkeypox virus using skin lesion images. They tested the dataset on five pre-trained deep neural networks: GoogLeNet, Places365-GoogLeNet, SqueezeNet, AlexNet, and ResNet-18. Transfer learning models such as Xception, DenseNet-169, and MobileNetV2 were also employed in other studies related to monkeypox diagnosis. Data augmentation techniques were applied to increase the size of the dataset and improve the performance of the deep neural networks. The deep learning models achieved high accuracy in detecting monkeypox virus from skin lesion images, with ResNet-18 obtaining the highest accuracy of 99.49%. Other performance metrics such as precision, recall, f1-score, and AUC were also considered, and the modified models obtained validation accuracies above 95%. Authors R33 employed deep convolutional neural networks of the residual network family, including ResNet-18, ResNet-50, and ResNet-101, to solve the vanishing gradient problem and achieve high performance in image classification. They also used SqueezeNet as a baseline model. The researchers achieved an average accuracy of 91.19% and an F1-score of 92.55% for the Monkeypox class using deep learning models. The researchers also used explainable artificial intelligence (XAI) techniques, such as LIME (Local Interpretable Model-Agnostic Explanations), to visualize the predictions made by the convolutional neural network. Authors R33 integrate deep neural networks with federated learning (FL) to address the challenges of medical data categorization. Also,they used a cycle-consistent generative adversarial network (GAN) to augment data samples for training. And utilize deep learning models such as MobileNetV2, Vision Transformer (ViT), and ResNet50 for the classification task. Authors R35 developed a hybrid artificial intelligence system using deep learning models (CSPDarkNet, InceptionV4, MnasNet, MobileNetV3, RepVGG, SE-ResNet, and Xception) and a long short-term memory (LSTM) model for monkeypox detection. The classification results were analyzed using metrics such as test accuracy and Cohen’s kappa score.The proposed hybrid artificial intelligence system achieved a test accuracy of 87% and a Cohen’s kappa score of 0.8222 for monkeypox detection. R36 presented a model, PoxNet22, by fine-tuning the InceptionV3 to classify monkeypox from other poxes. They used different preprocessing techniques to improve the image quality and used augmentation to reduce the possibility of overfitting. They got very high precession, recall and accuracy. R37 used transfer learning to assess the performance of different DL mehtods namely, VGG16, InceptionResNetV2, ResNet50, ResNet101, MobileNetV2, VGG19, and Vision Transformer (ViT) for diagnosing monkeypox. Among their models, modified VGG16 and MobileNetV2 outperformed others. For proving model’s transparency, they used LIME with their model. R38 presented a model for detecting monkeypox by using attention-based MobileNetV2. They used diversified performance parameters to asses the superiority of the proposed model. They also used Grad-CAM and LIME for ensuring transparency and interpretability of their model. Table 1 presents a comprehensive comparison of various parameters employed in previous studies and our study focused on monkeypox lesion detection. Each row corresponds to a specific study, while the columns represent different evaluation metrics that include accuracy, precision, recall, F1-score, area under the curve (AUC), etc. This analysis offers valuable insights into the existing literature and highlights the unique contribution of our study by incorporating all key evaluation metrics for a comprehensive evaluation of monkeypox lesion detection. Our study aims to contribute to the existing literature on monkeypox lesion detection by providing a comprehensive comparison of key evaluation metrics, including multiple training accuracy,multiple testing accuracy, Training loss, Testing Loss, precision, recall, F1-score, and area under the curve (AUC). As demonstrated in Table 1, our analysis revealed that several performance metrics were not consistently reported across existing studies. This observation underscores the need for a more standardized and comprehensive approach to evaluating monkeypox lesion detection models. By systematically comparing the performance of various pre-trained models and proposing a novel lightweight CNN model, MpoxSLDNet, our study addresses the limitations identified in previous research. Specifically, we have incorporated all key evaluation metrics, including multiple training accuracy,multiple testing accuracy, Training loss, Testing Loss, precision, recall, F1-score, and area under the curve (AUC), which were often overlooked in previous studies. This holistic evaluation approach provides a more robust assessment of model performance and facilitates a deeper understanding of the strengths and weaknesses of different monkeypox lesion detection methodologies.
@@
Table 1: Comparative analysis of parameters used in existing studies with our study.

### v2

@@ Abstract
Monkeypox virus (MPXV) is a zoonotic virus that poses a significant threat to public health, particularly in remote parts of Central and West Africa. Early detection of monkeypox lesions is crucial for effective treatment. However, due to its similarity with other skin diseases, monkeypox lesion detection is a challenging task. To detect monkeypox, many researchers used various deep-learning models such as MobileNetv2, VGG16, ResNet50, InceptionV3, DenseNet121, EfficientNetB3, MobileNetV2, and Xception. However, these models often require high storage space due to their large size. This study aims to improve the existing challenges by introducing a CNN model named MpoxSLDNet (Monkeypox Skin Lesion Detector Network) to facilitate early detection and categorization of Monkeypox lesions and Non-Monkeypox lesions in digital images. Our model represents a significant advancement in the field of monkeypox lesion detection by offering superior performance metrics, including precision, recall, F1-score, accuracy, and AUC, compared to traditional pre-trained models such as VGG16, ResNet50, and DenseNet121. The key novelty of our approach lies in MpoxSLDNet’s ability to achieve high detection accuracy while requiring significantly less storage space than existing models. By addressing the challenge of high storage requirements, MpoxSLDNet presents a practical solution for early detection and categorization of monkeypox lesions in resource-constrained healthcare settings. In this study, we have used "Monkeypox Skin Lesion Dataset" comprising 1428 skin images of monkeypox lesions and 1764 skin images of Non-Monkeypox lesions. Dataset’s limitations could potentially impact the model’s ability to generalize to unseen cases. However, the MpoxSLDNet model achieved a validation accuracy of 94.56%, compared to 86.25%, 84.38%, and 67.19% for VGG16, DenseNet121, and ResNet50, respectively. In addition, an average precision, recall, F1 score of 0.94 respectively, and AUC of 0.94 was recorded for our proposed model.
I Introduction
@@
Figure 1 provides an introductory overview of the efforts to enhance monkeypox lesion detection. It begins by elucidating the characteristics of the Monkeypox virus including its origins, classification, symptoms, and modes of transmission. Monkeypox virus (MPXV) is a zoonotic virus that is predominantly found in distant Central and West African regions . It was first identified in a monkey in 1959, when outbreaks of a disease resembling pox were observed in monkeys retained for research purposes, in a laboratory in Copenhagen, at a Danish research institute, and thus the name Monkeypox virus [2, 3]. It is a type of DNA virus classified within the orthopoxvirus genus and is a member of the family of Poxviridae, is closely linked to the Variola virus, which is the infectious illness that causes monkeypox [4]. In 1970, the first recorded instance of monkeypox cases in humans was documented in the Democratic Republic of the Congo, within an intensified campaign to wipe out smallpox. The monkeypox virus falls into two groups: Congo Basin (Central African) and West African. According to a previous report, the Congo Basin clade tends to produce severe disease more often, with case fatality ratios (CFR) of up to 10%. As of right now, the Democratic Republic of Congo reports a 3% Case Fatality Ratio among suspected cases. In the past, the West African clade has been attributed to a lower overall CFR of roughly 1% in an African population who are normally younger. An average incubation period is 12 days for MPXV, yet it might vary from 5 to 21 days. [5]. Monkeypox virus signs involve fever, headache, tenderness in the muscles, backaches, swelling of the lymph nodes, chills, as well as exhaustion [6, 7, 31]. The monkeypox virus circulates from a human to another human being thru droplets from the nose or throat, interaction with affected bodily fluids or contaminated surfaces, infectious rash, and sexual contact. This virus can also spread through the handling of infected animals [8].
There is unfortunately no medication or vaccination specifically against the monkeypox virus. Although supportive care can be given to combat symptoms and undesirable side effects. Antiviral drugs can be used that are used for treating other orthopoxvirus infections [9]. A skin lesion test utilizing the technique of electron microscopy or polymerase chain reaction, also known as a PCR test is time-consuming and takes a long time to detect pox [10]. The insufficient availability of PCR test kits, as well as the strict requirements for laboratory settings in research facilities, will delay the exact identification of suspects [11].
The detection of monkeypox lesions can be challenging due to the similarity that exists between monkeypox lesions and another pox [12, 13], which can lead to misdiagnosis. In addition, personnel in healthcare might have been unfamiliar with a clinical presentation of the lesions of monkeypox [14], which can further complicate accurate diagnosis. Moreover, physical examinations to diagnose monkeypox can increase the risk of transmission to healthcare professionals [15] and other patients. According to [16], isolation is recommended for a patient with a suspected monkeypox infection.
To address all these challenges, there is a need for the development of a reliable and accurate diagnostic tool that can quickly and non-invasively detect and classify monkeypox lesions from other poxvirus infections which also can significantly reduce the risk of transmission to healthcare professionals and other patients. Neural networks can help in this situation. With advances in the development involving AI models, many models based on deep learning have been utilized in the identification of various diseases [17, 18, 19, 30]. So, in our current research, we have developed a model named “MpoxSLDNet” based on convolutional neural networks (CNNs) to automatically detect monkeypox lesions from digitized pictures of human skin lesions.
@@
From the existing research work regarding monkeypox lesion detection, we can observe, in [1], authors used publicly available datasets. They presented an Android mobile application in which low modified deep learning model MobileNetv2 is used to classify the monkeypox lesion and non-monkeypox lesions with 91.11% accuracy. In [11], authors developed a monkeypox skin lesion dataset, and they employed deep learning networks including VGG16, ResNet50, InceptionV3, and Ensemble. In their study, among the evaluated models, ResNet50 demonstrated the highest overall accuracy, achieving an accuracy of 82.96%. Also VGG16, InceptionV3,Ensemble obtained accuracy of 81.48%, 74.07%,79.26% respectively. Here, the authors presented a web app that used the best-performing model. Apart from this, Authors [20] used publicly available datasets and employed five deep learning models, including DenseNet121, VGG19, MobileNetV2, EfficientNetB3, and Xception combined with the CBAM and Dense layers for categorization. Here, VGG19, DenseNet121, EfficientNetB3, MobileNetV2, and Xception achieved accuracy 71.86%, 78.27%, 81.43%, 74.07%, 83.89% respectively. Despite the significant efforts in the existing research work regarding monkeypox lesion detection, our model achieves superior results on the basis of accuracy, precision, recall, F1-score, and AUC compared to the state-of-the-art models. While the existing thesis work on monkeypox lesion detection has employed several deep-learning models, these models require significant storage space due to their large number of parameters and also require more computational time. In contrast, our proposed CNN model is designed to be both accurate and efficient, with fewer parameters that require a smaller amount of storage as well as computational time. This makes our approach more practical for use in resource-constrained settings, such as remote areas where monkeypox outbreaks are more likely to occur.
Our proposed MpoxSLDNet model is a Sequential neural network containing max pooling, convolutional layers, batch normalization, dropout, as well as fully connected layers. It takes input images of size 224x224x3 and learns to classify them as either having monkeypox lesions or not. The feature extraction process in our model involves using a series of convolutional layers of adding depth in addition to reducing spatial dimensions, commencing with max pooling layers to reduce the dimensionality of the feature maps. The extracted features are then flattened and passed through various fully connected layers with batch normalization and dropout regularization to stay away from overfitting. The final layer of output uses a sigmoid activation function to acquire a binary classification result. The contribution of this study to the field of monkeypox lesion detection is listed below:
@@ II Literature Review
Monkeypox is a rare and serious viral disease that presents a significant public health risk. The timely detection of monkeypox lesions is essential for effective treatment as well as outbreak prevention. Deep learning-based models have gained prominence in recent years due to their potential use in automating the detection of monkeypox lesions. A recent study conducted by [4] collected monkeypox images from various sources and developed a dataset of monkeypox images that are publicly available. This study established a modified VGG16 model to classify monkeypox patients and non-monkeypox patients. Post-image analysis explanations using a technique called Local Interpretable Model-Agnostic Explanations (LIME) were also used in order to verify outcomes. The modified VGG16 model was evaluated for two distinct studies, including study one (without augmentation) and study two (with augmentation). The achieved accuracies for studies one and two were 97 ± 1.8% and 88 ± 0.8%, respectively. However, this study only considered the modified VGG16 model. Additionally, Another study conducted by [21], proposed a model named MonkeypoxHybridNet based on CNN, which is comprised of VGG19, ResNet50, and InceptionV3. The performance of the model was assessed by evaluating its precision, F1 score, and accuracy against three pre-trained models, including VGG19, ResNet50, and InceptionV3. The achieved accuracy for VGG19, ResNet50, InceptionV3, and MonkeypoxHybridNet were 0.595, 0.705, 0.80, and 0.842, respectively. The proposed model showed significant results, but it has some limitations. The authors only presented accuracy, precision, F1 score, and confusion matrix but did not show the Recall and AUC score for the proposed MonkeypoxHybridNet model. In a subsequent study by [14], a variety of pre-trained models were used to distinguish between chickenpox and monkeypox lesions, including Inception V3, ResNet50, VGG16, VGG19, and AlexNet. They also implemented a CNN model. The chickenpox dataset used in the study was publicly available, while the monkeypox dataset was obtained through web scraping.VGG16 and VGG19 accuracy was also 80%, while ResNet50, AlexNet, and Inception V3 each provided accuracy results of 82.00%, 98.00%, and 89.00%, respectively. The recorded accuracy of the proposed CNN model was 99%. However, the AUC score, which is an important measure of classifier performance, was not calculated in this study. Apart from this, another study by [2] compared the performance of 13 pre-trained models to detect the monkeypox lesion. Universal custom layers had been used for all models, and the performance was analyzed based on four metrics: accuracy, precision, recall, and F1-score. The study identified DL models that exhibited the highest performance and ensembled them via majority voting on their probabilistic outputs to enhance the overall performance. The experiments utilized a dataset that is publicly accessible and the proposed ensemble method provided accuracy, precision, recall, and F1-score of 87.13%, 85.44%, 85.47%, and 85.40%, respectively. However, this study did not propose any lightweight CNN models, and the AUC score was not considered. In a subsequent study conducted by [11], A dataset called the Monkeypox Skin Lesion Dataset (MSLD) was created, which includes images of skin lesions associated with monkeypox, chickenpox, and measles. Three widely used pre-trained models, including InceptionV3, ResNet50, and VGG16, were utilized to identify monkeypox and non-monkeypox patients. Additionally, an ensemble of three models (InceptionV3, ResNet50, and VGG16) was employed in this study. The achieved accuracies were 81.48% (±6.87%), 82.96% (±4.57%), 74.07% (±3.78%), and 79.26% (±1.05%) for VGG16, ResNet50, InceptionV3, and the ensemble method, respectively. Furthermore, a web application was employed as an online monkeypox screening tool. It is known that pre-trained models require more storage space and computational time. In this work, only pre-trained models were utilized, and no lightweight CNN models were established. As for performance evaluation metrics, the AUC score was not considered. Furthermore, a study conducted by [22], used five pre-trained models including VGG16, VGG19, MobileNetV2, ResNet50, and EfficientNetB3. The performance of the model is assessed by measuring its accuracy, recall, precision, and F1 score. Their experimental result demonstrates that among the five pre-trained models MobileNetV2 gives the best classification performance. Achieved accuracy, precision, recall, and F1-score of MobileNetV2 are 98.16%, 0.99, 0.96, and 0.98. Their objective was to select the best-performing model to detect and classify monkeypox. In this study, the authors don’t consider the AUC score and don’t show the multiple training and testing accuracy. They also don’t discuss losses during training. Another study employed by [23], a newly curated image dataset, referred to as the "Monkeypox skin images dataset," was developed, encompassing four classes: monkeypox, chickenpox, measles, and normal images. In this study, five machine learning classifiers, including Logistic Regression, Random Forest, Support Vector Machine, K-Nearest Neighbor, as well as Extreme Gradient Boosting were used for classification. For feature extraction, VGG16, ResNet50, MobileNetV1, Inception V3, Xception. Also, they presented a novel CNN model called MonkeyNet, which is a modified version of DenseNet-201, and evaluated its performance. The performance of the model was assessed using various metrics including accuracy, precision, recall, F1-score, and AUC. Resulted in accuracy for the original dataset and augmented dataset of the proposed model are 91.91% and 98.91%, consecutively. Another study conducted by [20] utilized five deep learning models including VGG19, Xception, DenseNet121, MobileNetV2, and EfficientNetB3.To utilize a more relevant feature map, they integrate the deep learning model with the Convolutional Block Attention Module. This module encompasses mechanisms such as channel attention and spatial attention-based mechanism. Then dense layer, flattened layer, and classification layer are added to each model to classify monkeypox and others. The model’s performance is assessed using accuracy, precision, recall, and F1-score as evaluation metrics. Among the various combinations, Xception-CBAM-Dense yielded the highest performance with a validation accuracy of 83.89%, precision of 90.70%, recall of 89.10%, and F1-score of 90.11%. Authors [33] used deep learning techniques, specifically deep neural networks, for the detection of monkeypox virus using skin lesion images. They tested the dataset on five pre-trained deep neural networks: GoogLeNet, Places365-GoogLeNet, SqueezeNet, AlexNet, and ResNet-18. Transfer learning models such as Xception, DenseNet-169, and MobileNetV2 were also employed in other studies related to monkeypox diagnosis. Data augmentation techniques were applied to increase the size of the dataset and improve the performance of the deep neural networks. The deep learning models achieved high accuracy in detecting monkeypox virus from skin lesion images, with ResNet-18 obtaining the highest accuracy of 99.49%. Other performance metrics such as precision, recall, f1-score, and AUC were also considered, and the modified models obtained validation accuracies above 95%. Authors [34] employed deep convolutional neural networks of the residual network family, including ResNet-18, ResNet-50, and ResNet-101, to solve the vanishing gradient problem and achieve high performance in image classification. They also used SqueezeNet as a baseline model. The researchers achieved an average accuracy of 91.19% and an F1-score of 92.55% for the Monkeypox class using deep learning models. The researchers also used explainable artificial intelligence (XAI) techniques, such as LIME (Local Interpretable Model-Agnostic Explanations), to visualize the predictions made by the convolutional neural network. Authors [34] integrate deep neural networks with federated learning (FL) to address the challenges of medical data categorization. Also,they used a cycle-consistent generative adversarial network (GAN) to augment data samples for training. And utilize deep learning models such as MobileNetV2, Vision Transformer (ViT), and ResNet50 for the classification task. Authors [36] developed a hybrid artificial intelligence system using deep learning models (CSPDarkNet, InceptionV4, MnasNet, MobileNetV3, RepVGG, SE-ResNet, and Xception) and a long short-term memory (LSTM) model for monkeypox detection. The classification results were analyzed using metrics such as test accuracy and Cohen’s kappa score.The proposed hybrid artificial intelligence system achieved a test accuracy of 87% and a Cohen’s kappa score of 0.8222 for monkeypox detection. [37] presented a model, PoxNet22, by fine-tuning the InceptionV3 to classify monkeypox from other poxes. They used different preprocessing techniques to improve the image quality and used augmentation to reduce the possibility of overfitting. They got very high precession, recall and accuracy. [38] used transfer learning to assess the performance of different DL mehtods namely, VGG16, InceptionResNetV2, ResNet50, ResNet101, MobileNetV2, VGG19, and Vision Transformer (ViT) for diagnosing monkeypox. Among their models, modified VGG16 and MobileNetV2 outperformed others. For proving model’s transparency, they used LIME with their model. [39] presented a model for detecting monkeypox by using attention-based MobileNetV2. They used diversified performance parameters to asses the superiority of the proposed model. They also used Grad-CAM and LIME for ensuring transparency and interpretability of their model.
Table I presents a comprehensive comparison of various parameters employed in previous studies and our study focused on monkeypox lesion detection. Each row corresponds to a specific study, while the columns represent different evaluation metrics that include accuracy, precision, recall, F1-score, area under the curve (AUC), etc. This analysis offers valuable insights into the existing literature and highlights the unique contribution of our study by incorporating all key evaluation metrics for a comprehensive evaluation of monkeypox lesion detection. Our study aims to contribute to the existing literature on monkeypox lesion detection by providing a comprehensive comparison of key evaluation metrics, including multiple training accuracy,multiple testing accuracy, Training loss, Testing Loss, precision, recall, F1-score, and area under the curve (AUC). As demonstrated in Table I, our analysis revealed that several performance metrics were not consistently reported across existing studies. This observation underscores the need for a more standardized and comprehensive approach to evaluating monkeypox lesion detection models. By systematically comparing the performance of various pre-trained models and proposing a novel lightweight CNN model, MpoxSLDNet, our study addresses the limitations identified in previous research. Specifically, we have incorporated all key evaluation metrics, including multiple training accuracy,multiple testing accuracy, Training loss, Testing Loss, precision, recall, F1-score, and area under the curve (AUC), which were often overlooked in previous studies. This holistic evaluation approach provides a more robust assessment of model performance and facilitates a deeper understanding of the strengths and weaknesses of different monkeypox lesion detection methodologies.
@@
TABLE I: Comparative analysis of parameters used in existing studies with our study.